if you add5 to the sum of bubba's age and his poppa's age the total is 48. bubba's poppa is 5 years less than 3 times the age of bubba. how old is bubba and his poppa?
Answer by tutorcecilia(2152)

Bubb + Poppa=48 [sum of bubba's age and his poppa's age the total is 48]
Bubba=B
Poppa=3B-5 [poppa is 5 years less than 3 times the age of bubba]
.
B+3B-5=48 [Plug-in the values]
4B-5=48 [Solve for B]
4B-5+5=48+5
4B=53
4B/4=53/4
B=53/4
.
So, Bubba = 53/4=13.25 years old
Poppa = (3)(13.25)-5=34.75 years old
.
checking gives 13.25+34.75=48